# COBRA: A Continual Learning Approach to Vision-Brain Understanding

### v1

@@ 4.3 Rehearsal-based Continual Learning
| Method | Buffer | (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | (3,4,6,8)-(1,2,5,7) (2 steps) | ||||||||||
| (3,4,6,8) | (1,2,5,7) | all | (3,4,6,8) | (1,2,5,7) | all | ||||||||
| [4] | 500 | 0.283 | 82.5 | 0.317 | 83.6 | 0.300 | 83.1 | 0.286 | 83.9 | 0.319 | 86.2 | 0.303 | 85.1 |
| COBRA (Ours) | 0.315 | 93.3 | 0.352 | 93.8 | 0.334 | 93.6 | 0.317 | 93.7 | 0.354 | 94.3 | 0.336 | 94.0 | |
| [4] | 1500 | 0.292 | 85.7 | 0.324 | 86.4 | 0.308 | 86.1 | 0.295 | 87.2 | 0.327 | 90.3 | 0.311 | 88.8 |
| COBRA (Ours) | 0.322 | 94.9 | 0.357 | 94.5 | 0.340 | 94.7 | 0.322 | 95.2 | 0.358 | 95.2 | 0.340 | 95.2 | |
| [4] | 4000 | 0.302 | 88.9 | 0.335 | 89.1 | 0.319 | 89.0 | 0.308 | 89.8 | 0.338 | 92.8 | 0.323 | 91.3 |
| COBRA (Ours) | 0.334 | 95.6 | 0.362 | 95.9 | 0.343 | 95.8 | 0.326 | 96.3 | 0.363 | 96.6 | 0.345 | 96.5 | |
@@ 4.6 Ablation Studies
Effectiveness of the MRIFormer module. In COBRA, our MRIFormer includes a transformer encoder to accumulate subject-specific and individual information, and a transformer decoder to translate fMRI data into CLIP space. As shown in Table 5, compared to the linear layers used in prior studies, our module achieves approximately 3% higher performance in both SSIM and CLIP, which demonstrates the effectiveness of our approach.
Effectiveness of prompt length for PSS module. We evaluate the effectiveness of prompt length in the PSS module. In particular, we experiment with different top-k values as in the Eqn (3). The results are shown in Table 6. We found that selecting the top-k of 30 tokens for constructing the subject-specific prompt is the most optimal. Increasing this top-k number does not improve overall performance, but increases network size.
@@
Model sizes during continual learning. The size of COBRA grows gradually in response to an increasing number of subjects in CVBU. We simulate COBRA’s size as the number of subjects increases from 1 to 100. Additionally, we compare its model size to MindBridge[72] and MindEye2 [63], designed as a unified model. The results are shown in Fig. 6. Notably, the model size increases approximately nine times and three times for MindBridge and MindEye2 approximately as the subject count rises from 1 to 100. Meanwhile, COBRA’s size grows minimally.
This is because only the PSS and MRIFormer modules expanded to additional subjects. Designed as lightweight and simple modules, the PSS and MRIFormer module adds relatively few parameters compared to MindBridge and MindEye2.
@@ 5 Conclusions
This work has presented one of the first continual learning approaches to the Vision-Brain Understanding problems. In particular, we have introduced the new Subject Commonality (SC) module to capture vision-brain patterns shared across subjects. In addition, the Prompt-based Subject-Specific (PSS) module has been designed to learn unique vision-brain patterns for each subject. The MRIFormer is also incorporated to align fMRI features with consistent CLIP features. In this continual learning setup, only the PSS module needs updating to accommodate new subjects, effectively addressing catastrophic forgetting. Our COBRA approach has achieved state-of-the-art performance in the VBU tasks, as demonstrated on the NSD benchmark.
Limitations and Future Work. Our SC module may be limited when the number of subjects is only one, as it does not effectively capture patterns shared across multiple subjects. The more subjects the module is trained on, the more accurately it can identify common patterns. In addition, this module is currently limited to 80 classes of MS-COCO. These limitations will motivate new research in future work.
Broader Impact. The advancements in deep learning and neuroscience hold transformative potential for adaptability in neuroscience applications. Implementing continual learning will empower researchers to identify common fMRI features based on a given condition in the future. This will facilitate more adaptable and sophisticated applications involving fMRI, such as the reconstruction of visual perception, decoding of internal speech, and development of neural prosthetics for communication.
Supplementary Material
6 Supplementary

### v3

@@ 4.3.1 Rehearsal-based Vision-Brain Reconstruction Task
| (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | |||||||
| Method | Buffer | (3,4,6,8) | (1,2,5,7) | all | |||
| SSIM | CLIP | SSIM | CLIP | SSIM | CLIP | ||
| [41] | 500 | 0.2830.004 | 82.50.5 | 0.3170.004 | 83.60.5 | 0.3000.004 | 83.10.5 |
| COBRA | 0.3150.003 | 93.30.3 | 0.3520.003 | 93.80.3 | 0.3340.003 | 93.60.3 | |
| [41] | 1500 | 0.2920.004 | 85.70.5 | 0.3240.004 | 86.40.5 | 0.3080.004 | 86.10.5 |
| COBRA | 0.3220.003 | 94.90.2 | 0.3570.003 | 94.50.2 | 0.3400.003 | 94.70.2 | |
| [41] | 4000 | 0.3020.004 | 88.90.4 | 0.3350.004 | 89.10.4 | 0.3190.004 | 89.00.4 |
| COBRA | 0.3340.003 | 95.60.2 | 0.3620.003 | 95.90.2 | 0.3430.003 | 95.80.2 | |
| (3,4,6,8)-(1,2,5,7) (2 steps) | |||||||
| Method | Buffer | (3,4,6,8) | (1,2,5,7) | all | |||
| SSIM | CLIP | SSIM | CLIP | SSIM | CLIP | ||
| [41] | 500 | 0.2860.004 | 83.90.5 | 0.3190.004 | 86.20.5 | 0.3030.004 | 85.10.5 |
| COBRA | 0.3170.003 | 93.70.3 | 0.3540.003 | 94.30.3 | 0.3360.003 | 94.00.3 | |
| [41] | 1500 | 0.2950.004 | 87.20.5 | 0.3270.004 | 90.30.5 | 0.3110.004 | 88.80.5 |
| COBRA | 0.3220.003 | 95.20.2 | 0.3580.003 | 95.20.2 | 0.3400.003 | 95.20.2 | |
| [41] | 4000 | 0.3080.004 | 89.80.4 | 0.3380.004 | 92.80.4 | 0.3230.004 | 91.30.4 |
| COBRA | 0.3260.003 | 96.30.2 | 0.3630.003 | 96.60.2 | 0.3450.003 | 96.50.2 | |
4.3.2 Rehearsal-based Brain Captioning Task
Building on the models trained for the Rehearsal-based Vision-Brain Reconstruction task described in Section 4.3.1, similar to the brain captioning task of Section 4.2.2, we project the fMRI features embedding to language space using an adapter and generate the caption for the corresponding brain signal. We report the performance for this brain captioning task across multiple metrics: BLEU-1 and BLEU-2 in Table 9, BLEU-3 and BLEU-4 in Table 10, METEOR and ROUGE in Table 11, CIDEr and SPICE in Table 12, and CLIP-S and RefCLIP-S in Table 13. Generally, the performance improves gradually as buffer size increases from 500 to 4000, closing the gap to the upper bound. Compared to the previous method, [41], our proposed approach, COBRA, demonstrates efficiency in continual learning performance.
| (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | |||||||
| Method | Buffer | (3,4,6,8) | (1,2,5,7) | all | |||
| BLEU-1 | BLEU-2 | BLEU-1 | BLEU-2 | BLEU-1 | BLEU-2 | ||
| [41] | 500 | 55.10.3 | 36.40.3 | 56.20.5 | 37.30.4 | 55.60.4 | 36.90.3 |
| COBRA | 58.90.5 | 39.90.4 | 59.80.3 | 40.70.5 | 59.40.4 | 40.30.3 | |
| [41] | 1500 | 56.40.4 | 37.40.2 | 57.30.4 | 38.30.4 | 56.80.5 | 37.90.3 |
| COBRA | 59.60.4 | 40.70.5 | 60.60.3 | 41.50.2 | 60.10.4 | 41.10.3 | |
| [41] | 4000 | 57.60.3 | 38.60.3 | 58.50.4 | 39.40.2 | 58.00.3 | 39.00.3 |
| COBRA | 60.50.4 | 41.60.5 | 61.80.4 | 42.80.4 | 61.20.2 | 42.20.5 | |
| (3,4,6,8)-(1,2,5,7) (2 steps) | |||||||
| Method | Buffer | (3,4,6,8) | (1,2,5,7) | all | |||
| BLEU-1 | BLEU-2 | BLEU-1 | BLEU-2 | BLEU-1 | BLEU-2 | ||
| [41] | 500 | 56.30.5 | 37.10.3 | 57.40.3 | 38.00.4 | 56.90.3 | 37.60.2 |
| COBRA | 60.50.3 | 41.10.3 | 61.50.4 | 42.30.5 | 61.00.4 | 41.80.4 | |
| [41] | 1500 | 57.60.4 | 38.30.5 | 58.70.2 | 39.20.4 | 58.20.3 | 38.80.5 |
| COBRA | 61.20.3 | 41.90.2 | 62.30.5 | 42.70.3 | 61.80.2 | 42.30.3 | |
| [41] | 4000 | 58.80.5 | 39.30.4 | 59.90.4 | 40.20.2 | 59.30.3 | 39.80.4 |
| COBRA | 61.20.2 | 41.70.4 | 62.40.5 | 42.80.3 | 61.80.2 | 42.30.3 | |
| (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | |||||||
| Method | Buffer | (3,4,6,8) | (1,2,5,7) | all | |||
| BLEU-3 | BLEU-4 | BLEU-3 | BLEU-4 | BLEU-3 | BLEU-4 | ||
| [41] | 500 | 23.50.2 | 15.00.2 | 24.70.2 | 15.90.2 | 24.10.2 | 15.50.2 |
| COBRA | 26.30.2 | 16.60.2 | 27.40.2 | 17.30.2 | 26.90.2 | 17.00.2 | |
| [41] | 1500 | 24.40.2 | 15.70.2 | 25.50.2 | 16.50.2 | 25.00.2 | 16.10.2 |
| COBRA | 27.00.2 | 17.30.2 | 28.20.2 | 18.00.2 | 27.60.2 | 17.70.2 | |
| [41] | 4000 | 25.40.2 | 16.50.2 | 26.50.2 | 17.30.2 | 25.90.2 | 16.90.2 |
| COBRA | 27.80.2 | 18.00.2 | 29.00.2 | 18.80.2 | 28.40.2 | 18.40.2 | |
| (3,4,6,8)-(1,2,5,7) (2 steps) | |||||||
| Method | Buffer | (3,4,6,8) | (1,2,5,7) | all | |||
| BLEU-3 | BLEU-4 | BLEU-3 | BLEU-4 | BLEU-3 | BLEU-4 | ||
| [41] | 500 | 24.40.2 | 15.60.2 | 25.40.2 | 16.50.2 | 24.90.2 | 16.10.2 |
| COBRA | 27.60.2 | 17.40.2 | 28.50.2 | 18.20.2 | 28.10.2 | 17.80.2 | |
| [41] | 1500 | 25.60.2 | 16.20.2 | 26.70.2 | 17.10.2 | 26.10.2 | 16.70.2 |
| COBRA | 28.40.2 | 18.10.2 | 29.40.2 | 18.90.2 | 28.90.2 | 18.50.2 | |
| [41] | 4000 | 26.70.2 | 17.30.2 | 27.80.2 | 18.10.2 | 27.20.2 | 17.70.2 |
| COBRA | 29.30.2 | 18.60.2 | 30.40.2 | 19.50.2 | 29.90.2 | 19.10.2 | |
| (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | |||||||
| Method | Buffer | (3,4,6,8) | (1,2,5,7) | all | |||
| METEOR | ROUGE | METEOR | ROUGE | METEOR | ROUGE | ||
| [41] | 500 | 17.20.4 | 39.50.3 | 18.00.5 | 40.40.2 | 17.60.3 | 40.00.4 |
| COBRA | 19.00.3 | 42.20.4 | 19.90.5 | 43.00.3 | 19.40.2 | 42.60.3 | |
| [41] | 1500 | 17.90.2 | 40.20.5 | 18.70.4 | 41.10.3 | 18.30.5 | 40.70.3 |
| COBRA | 19.50.4 | 42.70.3 | 20.50.2 | 43.50.4 | 20.00.4 | 43.10.5 | |
| [41] | 4000 | 18.60.3 | 40.90.3 | 19.40.2 | 41.80.5 | 19.00.5 | 41.40.2 |
| COBRA | 20.00.2 | 43.30.4 | 21.10.4 | 44.10.3 | 20.50.4 | 43.70.3 | |
| (3,4,6,8)-(1,2,5,7) (2 steps) | |||||||
| Method | Buffer | (3,4,6,8) | (1,2,5,7) | all | |||
| METEOR | ROUGE | METEOR | ROUGE | METEOR | ROUGE | ||
| [41] | 500 | 17.70.5 | 39.80.2 | 18.60.3 | 40.70.4 | 18.20.2 | 40.30.3 |
| COBRA | 20.20.4 | 43.30.2 | 21.10.3 | 44.10.5 | 20.70.3 | 43.70.4 | |
| [41] | 1500 | 18.40.3 | 40.50.3 | 19.30.5 | 41.40.2 | 18.90.4 | 41.00.2 |
| COBRA | 20.70.2 | 43.80.4 | 21.60.5 | 44.60.3 | 21.20.3 | 44.20.3 | |
| [41] | 4000 | 19.10.2 | 41.20.5 | 20.00.3 | 42.10.4 | 19.60.4 | 41.70.3 |
| COBRA | 21.30.3 | 44.40.5 | 22.30.4 | 45.20.2 | 21.80.3 | 44.80.4 | |
| (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | |||||||
| Method | Buffer | (3,4,6,8) | (1,2,5,7) | all | |||
| CIDEr | SPICE | CIDEr | SPICE | CIDEr | SPICE | ||
| [41] | 500 | 50.00.4 | 11.80.3 | 51.00.5 | 12.20.3 | 50.50.4 | 12.00.4 |
| COBRA | 58.00.3 | 12.60.4 | 59.00.4 | 13.00.2 | 58.50.3 | 12.80.3 | |
| [41] | 1500 | 51.50.2 | 12.20.5 | 52.70.3 | 12.60.4 | 52.10.4 | 12.40.3 |
| COBRA | 59.50.5 | 13.00.2 | 60.70.3 | 13.30.4 | 60.10.3 | 13.20.5 | |
| [41] | 4000 | 52.80.3 | 12.60.3 | 54.00.4 | 13.00.2 | 53.40.5 | 12.80.4 |
| COBRA | 60.50.2 | 13.20.4 | 61.60.4 | 13.40.5 | 61.10.3 | 13.30.3 | |
| (3,4,6,8)-(1,2,5,7) (2 steps) | |||||||
| Method | Buffer | (3,4,6,8) | (1,2,5,7) | all | |||
| CIDEr | SPICE | CIDEr | SPICE | CIDEr | SPICE | ||
| Method | Buffer | C (3,4,6,8) | S (3,4,6,8) | C (1,2,5,7) | S (1,2,5,7) | C (All) | S (All) |
| [41] | 500 | 51.50.5 | 12.50.2 | 52.60.3 | 12.80.3 | 52.10.2 | 12.60.4 |
| COBRA | 59.50.4 | 13.10.3 | 60.60.3 | 13.40.4 | 60.00.5 | 13.30.3 | |
| [41] | 1500 | 53.10.4 | 12.90.3 | 54.40.2 | 13.30.5 | 53.80.3 | 13.10.2 |
| COBRA | 61.30.3 | 13.30.2 | 62.40.5 | 13.50.4 | 61.80.3 | 13.40.4 | |
| [41] | 4000 | 54.60.3 | 13.20.4 | 55.90.4 | 13.40.2 | 55.20.5 | 13.30.3 |
| COBRA | 62.00.2 | 13.40.3 | 63.10.4 | 13.50.2 | 62.50.4 | 13.40.4 | |
| (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | |||||||
| Method | Buffer | (3,4,6,8) | (1,2,5,7) | all | |||
| CLIPS | RefCLIPS | CLIPS | RefCLIPS | CLIPS | RefCLIPS | ||
| [41] | 500 | 64.30.3 | 67.50.4 | 65.60.4 | 68.90.3 | 65.00.5 | 68.20.2 |
| COBRA | 66.00.2 | 70.00.5 | 67.30.4 | 71.50.3 | 66.60.4 | 70.80.4 | |
| [41] | 1500 | 65.20.4 | 68.60.3 | 66.50.5 | 70.10.2 | 65.90.2 | 69.30.3 |
| COBRA | 67.00.5 | 71.00.4 | 68.50.2 | 72.50.3 | 67.80.3 | 71.80.5 | |
| [41] | 4000 | 66.10.3 | 69.50.5 | 67.40.2 | 70.90.3 | 66.80.4 | 70.20.2 |
| COBRA | 68.10.4 | 72.10.3 | 69.40.3 | 73.40.2 | 68.70.5 | 72.70.3 | |
| (3,4,6,8)-(1,2,5,7) (2 steps) | |||||||
| Method | Buffer | (3,4,6,8) | (1,2,5,7) | all | |||
| CLIPS | RefCLIPS | CLIPS | RefCLIPS | CLIPS | RefCLIPS | ||
| [41] | 500 | 63.40.3 | 69.20.2 | 64.80.4 | 70.30.5 | 64.10.3 | 69.80.4 |
| COBRA | 68.00.4 | 73.30.4 | 69.50.3 | 74.00.3 | 68.80.2 | 73.70.5 | |
| [41] | 1500 | 64.70.4 | 70.80.5 | 66.00.2 | 71.80.3 | 65.40.3 | 71.30.2 |
| COBRA | 68.50.3 | 73.50.4 | 69.70.5 | 74.10.2 | 69.10.4 | 73.80.3 | |
| [41] | 4000 | 65.40.2 | 71.30.4 | 66.80.5 | 72.50.4 | 66.10.3 | 71.90.3 |
| COBRA | 69.90.4 | 73.40.2 | 71.40.3 | 74.70.5 | 70.70.3 | 74.10.2 | |
@@ 4.5 Comparison To Previous SOTA Models Of Vision-Brain Reconstruction
4.6 Comparison To Previous SOTA Models of Brain Captioning
Following the experimental setup and evaluation protocol from [26], we train our proposed COBRA model for the brain captioning task. As shown in Table 15, COBRA consistently achieves state-of-the-art results across all standard captioning metrics, including BLEU-1 to BLEU-4, METEOR, ROUGE, CIDEr, SPICE, CLIP-S, and RefCLIP-S. Compared to prior approaches such as SDRecon [87], OneLLM [88], BrainCap [89], and UMBRAE [26], COBRA demonstrates substantial improvements, particularly in CIDEr and CLIP-S metrics, which are indicative of higher semantic alignment with the ground truth captions. Notably, COBRA outperforms UMBRAE, which is the current SOTA, by a clear margin in CIDEr (63.1 vs. 61.06) and CLIP-S (69.2 vs. 67.78), while also achieving the best scores on BLEU-4 and METEOR. This suggests that COBRA not only captures low-level lexical cues but also preserves the high-level semantics of the original stimuli more effectively .Fig. 6 shows superior brain captioning results compared to prior work.
| Method | BLEU1 | BLEU2 | BLEU3 | BLEU4 | METEOR | ROUGE | CIDEr | SPICE | CLIP-S | RefCLIP-S |
| Shikra-w/img [83] | 82.38 | 69.90 | 58.63 | 49.66 | 35.60 | 65.49 | 161.43 | 27.62 | 80.60 | 85.92 |
| SDRecon [87] | 36.21 | 17.11 | 7.72 | 3.43 | 10.03 | 25.13 | 13.83 | 5.02 | 61.07 | 66.36 |
| OneLLM [88] | 47.04 | 26.97 | 15.49 | 9.51 | 13.55 | 35.05 | 22.99 | 6.26 | 54.80 | 61.28 |
| UniBrain [90] | - | - | - | - | 16.90 | 22.20 | - | - | - | - |
| BrainCap [89] | 55.96 | 36.21 | 22.70 | 14.51 | 16.68 | 40.69 | 41.30 | 9.06 | 64.31 | 69.90 |
| UMBRAE-S1 [26] | 57.63 | 38.02 | 25.00 | 16.76 | 18.41 | 42.15 | 51.93 | 11.83 | 66.44 | 72.12 |
| UMBRAE [26] | 59.44 | 40.48 | 27.66 | 19.03 | 19.45 | 43.71 | 61.06 | 12.79 | 67.78 | 73.54 |
| COBRA | 61.2 | 42.4 | 30.1 | 19.4 | 21.7 | 44.3 | 63.1 | 12.6 | 69.2 | 73.5 |
@@ 4.7 Ablation Studies
Effectiveness of the MRIFormer module. In COBRA, our MRIFormer includes a transformer encoder to accumulate subject-specific and individual information, and a transformer decoder to translate fMRI data into CLIP space. As shown in Table 17, our module outperforms the linear layers used in prior studies by approximately 3% in both SSIM and CLIP, demonstrating the effectiveness of our approach.
Effectiveness of prompt length for PSS module. We evaluate the effectiveness of prompt length in the PSS module. In particular, we experiment with different top- values as in the Eqn (4). The results are shown in Table 18. We observed that consistently yields the best performance in both SSIM and CLIP metrics. Larger did not improve performance and increased model size; smaller resulted in underfitting the subject-specific space. Thus, balances expressiveness and efficiency.
@@
Model sizes during continual learning. The size of COBRA grows gradually in response to an increasing number of subjects in CVBU. We simulate COBRA’s size as the number of subjects increases from 1 to 100. Additionally, we compare its model size to that of MindBridge [1] and MindEye2 [8], which are designed as unified models. The results are shown in Fig. 7. Notably, the model size increases approximately nine times and three times for MindBridge and MindEye2, respectively, as the subject count rises from 1 to 100. Meanwhile, COBRA’s size grows minimally. This is because only the PSS and MRIFormer modules were expanded to additional subjects. Designed as lightweight and simple modules, the PSS and MRIFormer module adds relatively few parameters compared to MindBridge and MindEye2.
Accuracy of SC and PSS Modules In the proposed COBRA, the SC module is designed to remain fixed when training on new subjects, as it captures common vision-brain patterns learned from previous subjects. To verify this, we trained the SC module on subjects (3,4,6,8) and tested it on the unseen subjects (1,2,5,7). The SC module achieved an F1 score of 0.84, supporting our hypothesis. For the PSS module, we evaluated its accuracy in learning subject-specific features for each subject. By using a prompt-based approach, the PSS achieved a high accuracy of 98.5%, indicating that the module effectively selects prompts that represent subject-specific features. Furthermore, we also design the PSS module following attention pooling [78] or random-based approaches to select subject-specific tokens. The performance of these approaches is 82.8% and 21.6%, respectively. These results demonstrate the efficiency of our PSS module using a prompt-based approach. The detailed results are presented in Table 16.
| Module | Approach | Metric | Accuracy |
| SC | ViT | f1 score | 0.84 |
| PSS | Prompt-based | accuracy | 98.5 |
| PSS | Attention Pooling [78] | accuracy | 82.8 |
| PSS | Random | accuracy | 21.6 |
Visualization of Subject-Specific Features In this section, we analyze the discrimination of subject-specific features, i.e., , that is extracted from the PSS module. We employ T-SNE [91] to reduce the dimension of the features. The visualization is illustrated in Fig. 8. The result demonstrates that the PSS module was able to learn the specific features that are discriminative for the subjects.
Qualitative Comparison With Previous Continual Learning Approaches. In this section, we compare the qualitative vision-brain reconstruction results of the model under four scenarios: without any continual learning (CL) approach, using LwF [82], using PLOP [81], and our proposed method, COBRA. The results are illustrated in Fig. 9 and Fig. 10. Our proposed method demonstrates significantly better and clearer reconstruction results compared to previous methods. In contrast, LwF [82] and PLOP [81] lose critical contextual details. Notably, without applying any CL approach, the generated images are incoherent and fail to reconstruct the context meaningfully.
Effect of Loss Balancing Hyperparameters. We conducted an ablation study to evaluate the effect of loss weighting, with a focus on the regularization loss , which prevents prompt center collapse in the PSS module. As shown in Table 19, when is removed (), SSIM drops dramatically from 0.329 to 0.257 (a relative drop of over 21%), and CLIP score falls from 92.9 to 88.4. This substantial degradation confirms that without sufficient regularization, the PSS prompt representations collapse to trivial solutions and fail to capture subject-specific variations. In contrast, performance is less sensitive to variations in , which controls alignment with CLIP features. These results highlight the critical importance of in maintaining representation diversity and preventing catastrophic collapse, justifying our default configuration of in all experiments.
@@ 5 Conclusions
This work has presented one of the first continual learning approaches to the Vision-Brain Understanding problems. In particular, we have introduced the new Subject Commonality (SC) module to capture vision-brain patterns shared across subjects. Additionally, the Prompt-based Subject-Specific (PSS) module has been designed to learn unique vision-brain patterns for each subject. The MRIFormer is also incorporated to align fMRI features with consistent CLIP features. In this continual learning setup, only the PSS module needs updating to accommodate new subjects, effectively addressing catastrophic forgetting. Our COBRA approach has achieved state-of-the-art performance in the VBU tasks, as demonstrated on the NSD benchmark.
Limitations and Future Work. Our SC module may be limited when the number of subjects is only one, as it does not effectively capture patterns shared across multiple subjects. The more subjects the module is trained on, the more accurately it can identify common patterns. In addition, this module is currently limited to 80 classes of MS-COCO. These limitations will motivate future research.
Broader Impact. Advancements in deep learning and neuroscience hold transformative potential for enhancing adaptability in neuroscience applications. Implementing continual learning will empower researchers to identify common fMRI features based on a given condition in the future. It will facilitate more adaptable and sophisticated applications involving fMRI, such as the reconstruction of visual perception, decoding of internal speech, and development of neural prosthetics for communication.
6 Data Availability
The data that support the findings of this study are openly available at https://naturalscenesdataset.org/.
7 Appendix
7.1 Explainations of Regions of Interest
Regions of Interest (ROIs) are specific brain areas responsible for various perceptual functions in humans. For example, and are responsible for processing face and body while we are looking at someone. It means that the neurons in these regions are firing, indicating that the stimulus actively activates them. These regions are determined by experiments with statistics. For instance, consider the floc-faces example. A set of visual stimuli is prepared to determine these regions, alternating between categories such as faces, objects, scenes, and words. fMRI signals are recorded from multiple participants as they are exposed to these stimuli. To pinpoint the floc-faces regions, the brain’s response to faces is contrasted with its responses to other categories (e.g., faces vs. objects) using statistical methods like General Linear Models (GLM).